OpenStudy (anonymous):

Which is a factor of 12x2 – 34x + 20 ?

OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 6 = x2 − x?

OpenStudy (anonymous):

there.isn't.any

OpenStudy (anonymous):

there isnt ,

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

how do yo get people to answer questions.. lol

OpenStudy (anonymous):

lol..idk

OpenStudy (anonymous):

4 6x – 5 x + 2 3x – 5

OpenStudy (anonymous):

factor the gcf

OpenStudy (anonymous):

which.is.2?

OpenStudy (anonymous):

2(6x^2-17+12)

OpenStudy (anonymous):

ok take common 2 from the given you get 2(6x^-17x+10) now break the midterm 2(6x^2-12x-5x+10) are you following?

OpenStudy (anonymous):

@TiaW

OpenStudy (anonymous):

@huneya. where.the.12.comefrom?

OpenStudy (anonymous):

yes.i.think.so

OpenStudy (anonymous):

do.u.break.it.down.again?

OpenStudy (anonymous):

ok (-12x-5x)=-17x i hope it is clear so now factor as 2((6x(x-2)-5(x-2)) are you getting this ??

OpenStudy (anonymous):

no...what.happen.to.the.17

OpenStudy (anonymous):

i just broke it as we can write -17x x as -12x-5x correct??

OpenStudy (anonymous):

@TiaW

OpenStudy (anonymous):

ok...i.think.i.got.it

OpenStudy (anonymous):

ok did you get this step? so now factor as 2((6x(x-2)-5(x-2))

OpenStudy (anonymous):

not.at.all

OpenStudy (anonymous):

@sami-21 i got 2(2x-3)(3x-4)

OpenStudy (anonymous):

ok 2(6x^2-12x-5x+10) taking 6x common from the first two terms from 6x^2-12x we can write as 6x(x-2) now taking -5 common from the last two terms from -5x+10 we can write it as -5(x-2) so combing the above two 2((6x(x-2)-5(x-2)) is it ok now ?

OpenStudy (anonymous):

@TiaW ,@huneya got it ?

OpenStudy (anonymous):

ok...i.got.it

OpenStudy (anonymous):

6x-5?

OpenStudy (anonymous):

is it okay now ?? can you do the rest by yourself just take x-2 common and there you are.

OpenStudy (anonymous):

@huneya is it ok?

OpenStudy (anonymous):

yeah.its.ok..thnxx

OpenStudy (anonymous):

yw:)